OpenStudy (anonymous):

Which is an equation in point-slope form for the given point and slope?Point: (7, 4); Slope: 6

OpenStudy (anonymous):

A. y – 7 = 6(x + 4) B. y – 4 = 6(x – 7) C. y – 4 = 6(x + 7) D. y + 4 = 6(x – 7)

OpenStudy (anonymous):

use slope point form we have point (x1 , y1)= (7 ,4 ) & slope m = 6 so we get y - y1 = m*(x - x1) just plug the values so we get y - 4 = 6 * (x - 7) so just try to simplify this equation ..

OpenStudy (anonymous):

Is it clear ?

OpenStudy (anonymous):

y – 4 = 6(x – 7)

OpenStudy (anonymous):

B